10 = 1/5 a + 2 solve the equation for a?
Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53750)

.
10 = 1/5 a + 2 solve the equation for a?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Step 1.  Subtract 2 from both sides.  You will get  

         10-2 = 1/5 a,  or  8 = 1/5 a.



Step 2.  Multiply both sides of the last equation by 5.  You will get 

             8*5 = a,  or  a = 40.


So, a = 40 is the solution and your ANSWER.



CHECK.  With a = 40, right side of the given equation is  1/5 40 + 2 = 8 + 2 = 10.

        So, right side is equal to the left side.

        It means that the equation is solved correctly.

My congrats !


The answer by @lynnlo is incorrect.